# Auditing a Dutch Public Sector Risk Profiling Algorithm Using an Unsupervised Bias Detection Tool

### v1

@@ 4 Detecting Bias in DUO’s Risk Profiling Algorithm
The unsupervised bias detection tool will be applied to the DUO dataset. Because this dataset includes ground truth labels for the demographic group, in the form of aggregated statistics on the migration background of students, our case study is suited to evaluate the effectiveness of the tool in detecting bias.
For the DUO audit, researchers were granted access to aggregated statistics on students’ migration by Statistics Netherlands [52]. If these data had not been available, the unsupervised bias detection tool discussed in this paper could have been used as an alternative methods for examining potential bias in the CUB process.
4.1 Description of the Dataset
@@
DUO’s risk profiling algorithm differentiated on the basis of three variables. First, the type of education the student is following, categorized as MBO 1-2, MBO 3-4 (vocational education) HBO (higher professional education) and WO (university education). Second, the age of the student, categorized as: 15-18, 19-20, 21-22, 23-24 and 25-50 years. Third, the distance between the address of the parents and the student, measured through the 6-digit postal code of the registered addresses and categorized as: 0km9990km indicates that student and parents share the same postal code but not the same house number., 1m-1km, 1-2km, 2-5km, 5-10km, 10-20km, 20-50km, 50-500km, unknown.
@@ 4.2 Applying the Unsupervised Biased Detection Tool
Data preprocessing: We transform the 3 categorical profiling characteristics to a set of 17 binary (non-overlapping) variables. We define no ordinal relationship between categories. We use the $k$-modes clustering within Algorithm 1.
Bias metric: We select whether a student is deemed “high risk” by the risk profiling algorithm or not as the bias metric. If the likelihood of being classified as “high risk” is equal for students of Dutch origin and those with a non-European migration background, it would correspond to the fairness metric— demographic parity [22]. Alternative definitions of bias often whether a college grant was duly granted ($y_{i}$), which in this case is only available for the small set of students who have manually been controlled by DUO ($n=3,179$). As classifications by the risk profiling algorithm contributed to discrimination in the CUB process [5], we consider this an appropriate bias metric. We emphasize that the HBAC algorithm can be used for each bias metric of interest.
@@ 4.3 Analyzing Identified Clusters
Figure 3: The percentage of students in each cluster for the three characteristics used in the risk profiling algorithm: (a) type of education, (b) age and (c) distance to parents within the student population of 2014, excluding students for which the risk profile was unknown ($n=214,599$). For each subgroup, the average is indicated by the dashed line.
@@ 4.4 Comparing Results with Aggregate Statistics on non-European Migration Background
Figure 4: The percentage of students with a non-European migration background for the three used profiling characteristics: (a) type of education, (b) age and (c) distance to parents for students in the CUB 2014 dataset ($n=248,650$), [^13] Since the aggregated data of Statistics Netherlands is rounded to the nearest 10, we report here $n=248,650$ instead of the $248,649$ for the CUB data. excluding students where the distance is unknown.
@@
Figure 5: The percentage of students with a non-European migration background for three (bivariate) combinations of the used profiling characteristics: (a) type of education and age, (b) type of education and distance to parents, and (c) age and distance to parents for the CUB 2014 dataset ($n=248,650$), [^15] Supra note 10excluding students where the distance is unknown.
@@ 6 Open Source Tool
The unsupervised bias detection tool is available on PyPI  [^16] https://pypi.org/project/unsupervised-bias-detection/. It implements the HBAC algorithm to detect potentially discriminated groups of users. Currently, the library supports k-means and k-modes algorithms for splitting the data into clusters. The modular design of the tool ensures that additional splitting methods can be integrated with minimal effort. To enhance accessibility, the tool is embedded within a web application, enabling people with no programming background to utilize its capabilities via an intuitive graphical interface [^17] https://algorithmaudit.eu/technical-tools/bdt/.
@@ 8 Conclusion
By advancing unsupervised learning methods, we enhance public knowledge on auditing algorithmic-supported decision-making processes on bias, in the absence of demographic data. The tool provides a scalable method that helps private entities, as well as internal and external auditors, to obtain inferences about seemingly neutral rules or practices that might embed unwanted biases. The tool could support further investigation into whether demographic groups face disadvantages from algorithmic-supported decision-making systems.
Through a simulation study, we highlight important considerations when applying unsupervised bias detection methods to avoid drawing incorrect conclusions. The simulations illustrate the necessity of using sample-splitting and Bonferroni corrections for valid inferences. The unsupervised bias detection method has been implemented as an open-source tool in the form of a Python package, accompanied by a graphical interface in a web application to encourage adoption by a non-technical audience. We hope that the contributions from this paper will serve the community by assisting in the detection of biases in algorithmic-supported decision-making systems.

### v2

@@ 4 Detecting Bias in DUO’s Risk Profiling Algorithm
The unsupervised bias detection tool will be applied to the DUO dataset. As requested from Statistics Netherlands, researchers were granted access to ground truth labels for demographic groups in the form of aggregated statistics on the migration background of students [52]. This case study is therefore well-suited to evaluate the effectiveness of the unsupervised tool in detecting bias. If these ground-truth group-level data had not been available, the unsupervised bias detection tool discussed in our paper could have served as method to examine potential bias in the CUB process.
Figure 3: The percentage of students in each cluster for the three characteristics used in the risk profiling algorithm: (a) type of education, (b) age, and (c) distance to parents within the student population of 2014, excluding students for which the risk profile was unknown ($n=214,599$). For each subgroup, the average is indicated by the dashed line.
4.1 Description of the DUO Dataset
@@
DUO’s risk profiling algorithm differentiated on the basis of three variables. First, the type of education the student is following, categorized as MBO 1-2, MBO 3-4 (vocational education) HBO (higher professional education) and WO (university education). Second, the age of the student, categorized as: 15-18, 19-20, 21-22, 23-24 and 25-50 years. Third, the distance between the address of the parents and the student, measured through the 6-digit postal code of the registered addresses and categorized as: 0km1010100km indicates that student and parents share the same postal code but not the same house number., 1m-1km, 1-2km, 2-5km, 5-10km, 10-20km, 20-50km, 50-500km, unknown.
@@ 4.2 Applying the Unsupervised Biased Detection Tool
Data preprocessing: We transform the 3 categorical profiling characteristics to one-hot encoded variables. We define no ordinal relationship between categories. We use the $k$-modes clustering within Algorithm 1.
Bias metric: As stated before, our metric requires that the probability of being classified as “high risk” by the risk profiling algorithm (based on $f(x_{i})$) is equal for students with a Dutch/European migration background and students with a non-European migration background. Thus, if these probabilities are equal, this condition satisfies our bias metric and corresponds to demographic parity [22]. We consider this an appropriate bias metric because classifications by the risk profiling algorithm contributed to discrimination in the CUB process [5]. In addition, the assessment of the adverse effect of the risk profiling algorithm on different demographic groups is a crucial aspect to investigate indirect discrimination under European non-discrimination law. An alternative definition of bias could involve examining whether college grants were duly awarded ($y_{i}$). However, ground truth labels indicating actual outcomes are only available for a small subset of students who underwent manual review by DUO (n=3,179). Consequently, we use predicted risk scores ($f(x_{i})$) instead, enabling the application to a substantially larger population (n=298,882). We emphasize that the HBAC algorithm can be used for any bias metric of interest.
@@ 4.3 Analyzing Identified Clusters
Figure 4: The percentage of students with a non-European migration background for the three used profiling characteristics: (a) type of education, (b) age and (c) distance to parents for students in the CUB 2014 dataset ($n=248,650$), excluding students where the distance is unknown.
@@ 4.4 Comparing Results with Aggregate Statistics on non-European Migration Background
Figure 5: The percentage of students with a non-European migration background for three (bivariate) combinations of the used profiling characteristics: (a) type of education and age, (b) type of education and distance to parents, and (c) age and distance to parents for the CUB 2014 dataset ($n=248,650$), excluding students where the distance is unknown.
@@ 6 Open Source Tool
The unsupervised bias detection tool is available on PyPI 131313link removed for anonymous review. It implements the HBAC algorithm for detecting the potentially discriminated groups of users. Currently, the library supports k-means and k-modes algorithms for splitting the data into clusters. The modular design of the tool ensures that additional splitting methods can be integrated with minimal effort. To enhance accessibility, the tool is embedded within a web application, enabling people with no programming background to utilize its capabilities via an intuitive graphical interface141414link removed for anonymous review.
@@ 8 Conclusion
By improving unsupervised learning methods, we enhance public knowledge on auditing algorithmic-supported decision-making processes on bias, in the absence of demographic data. The tool provides a scalable method that helps private entities, as well as internal and external auditors, to obtain inferences about seemingly neutral rules or practices that might embed unwanted biases. The tool could support further investigation into whether demographic groups face disadvantages from algorithmic-supported decision-making systems.
Through a simulation study, we highlight important considerations when applying unsupervised bias detection methods to avoid drawing incorrect conclusions. The simulations illustrate the necessity of using sample-splitting and Bonferroni corrections for valid inferences. In addition, the unsupervised bias detection method has been implemented as an open-source tool in the form of a Python package, accompanied by a graphical interface in a web application to encourage adoption by a non-technical audience. We hope that the contributions from this paper will serve the community by assisting in the detection of biases in algorithmic-supported decision-making systems.